Exercise 15-02

Mountain Corporation was organized on January 1, 2020. It is authorized to issue 9,500 shares of 8%, $100 par value preferred stock, and 523,200 shares of no-par common stock with a stated value of $1 per share. The following stock transactions were completed during the first year.

Jan. 10 Issued 80,270 shares of common stock for cash at $7 per share.
Mar. 1 Issued 5,750 shares of preferred stock for cash at $111 per share.
Apr. 1 Issued 24,580 shares of common stock for land. The asking price of the land was $91,860; the fair value of the land was $80,270.
May 1 Issued 80,270 shares of common stock for cash at $9 per share.
Aug. 1 Issued 9,500 shares of common stock to attorneys in payment of their bill of $51,700 for services rendered in helping the company organize.
Sept. 1 Issued 9,500 shares of common stock for cash at $11 per share.
Nov. 1 Issued 920 shares of preferred stock for cash at $106 per share.


Prepare the journal entries to record the above transactions. (Credit account titles are automatically indented when amount is entered. Do not indent manually. If no entry is required, select "No Entry" for the account titles and enter 0 for the amounts.)
Answer #1
Date Account Titles and Explanation Debit Credit
Jan. 10 Cash 561890 =80270*7
     Common stock 80270
     Paid-in Capital in Excess of Stated Value-Common Stock 481620
Mar. 1 Cash 638250 =5750*111
     Preferred stock 575000 =5750*100
     Paid-in Capital in Excess of Par value-Preferred Stock 63250
Apr. 1 Land 80270
     Common stock 24580
     Paid-in Capital in Excess of Stated Value-Common Stock 55690
May 1 Cash 722430 =80270*9
     Common stock 80270
     Paid-in Capital in Excess of Stated Value-Common Stock 642160
Aug. 1 Organization expense 51700
     Common stock 9500
     Paid-in Capital in Excess of Stated Value-Common Stock 42200
Sept. 1 Cash 104500 =9500*11
     Common stock 9500
     Paid-in Capital in Excess of Stated Value-Common Stock 95000
Nov. 1 Cash 97520 =920*106
     Preferred stock 92000 =920*100
     Paid-in Capital in Excess of Par value-Preferred Stock 5520